# Hadamard Representations: Augmenting Hyperbolic Tangents in RL

### v2

@@ Abstract
Activation functions are one of the key components of a deep neural network. The most commonly used activation functions can be classed into the category of continuously differentiable (e.g. tanh) and linear-unit functions (e.g. ReLU), both having their own strengths and drawbacks with respect to downstream performance and representation capacity through learning (e.g. measured by the number of dead neurons and the effective rank). In reinforcement learning, the performance of continuously differentiable activations often falls short as compared to linear-unit functions. We provide insights into the vanishing gradients associated with the former, and show that the dying neuron problem is not exclusive to ReLU’s. To alleviate vanishing gradients and the resulting dying neuron problem occurring with continuously differentiable activations, we propose a Hadamard representation. Using deep Q-networks and proximal policy optimization in the Atari domain, we show faster learning, a reduction in dead neurons and increased effective rank.
@@ 1 Introduction
In reinforcement learning (RL) (Sutton & Barto, 2018), even though training results in a large number of dying ReLU’s (Gulcehre et al., 2022; Sokar et al., 2023), the ReLU function still remains the most popular activation for performance reasons (Henderson et al., 2018). Similar to supervised learning (Teney et al., 2024), continuously differentiable activation functions such as the hyperbolic tangent are therefore not commonly used in RL (see Fig. 1). However, one might argue that their symmetrical, bounded shape and smooth gradient landscape offer optimization advantages that the ReLU lacks. Recent findings also indicate that a hidden layer activated by a hyperbolic tangent displays a high effective rank and thus a high layer expressivity (Kumar et al., 2021; Gulcehre et al., 2022). Despite being a theoretically sound candidate, its lack of success in RL has not been thoroughly investigated.
We therefore aim to provide insights into the hyperbolic tangent’s suboptimality, revealing that the vanishing gradient problem leads to dying neurons and under-utilization of the full network capacity. Based on these insights, we mitigate said effects by augmenting the original architecture. Specifically, we focus our research on the activations of the compressed representation of an encoder, defined as the final hidden layer, and provide an alternative to the conventional parameterization of this representation. Our contributions can be summarized as follows:
  • We show that, in reinforcement learning, dying hyperbolic tangents are a phenomenon of a similar scale as the dying ReLU problem.
@@
  • We empirically show that, without hyperparameter tuning or the use of auxiliary losses, a Hadamard representation yields significant performance gains over a standard hyperbolic tangent in the Atari domain, and reveal how it decreases dying neurons and increases the internal representations’ effective rank.
@@ 3 Augmenting Hyperbolic Tangents
Continuously differentiable activations such as the hyperbolic tangent ($tanh$) and the sigmoid ($\sigma$) activations are fundamentally different than the $ReLU$ or its Linear-Unit descendants, which are non-symmetric and have a large part of the input space mapped to zero (leading to sparsity). The hyperbolic tangent and the sigmoid output values in the ranges $[-1,1]$ and $[0,1]$, respectively. These functions are defined as:
$tanh(x)=\frac{e^{x}-e^{-x}}{e^{x}+e^{-x}},\qquad\sigma(x)=\frac{1}{1+e^{-x}}$ (1)
@@
Although common literature has focused on the dying ReLU problem (He et al., 2015; Lu et al., 2019; Gulcehre et al., 2022; Sokar et al., 2023), we find that hidden layers activated by hyperbolic tangents similarly show strong dying neuron behavior. When using a hyperbolic tangent in a deep neural network, a single neuron $\alpha_{i},i\in\mathbb{R}^{w}$, with $w$ the layer dimension, can be classified as dying if:
$|\alpha_{i}|\approx 1\;\forall\;s_{t}\in B$ (2)
Where $s_{t}$ is an observation in buffer $B$. Note that an absolute value is used, as a hyperbolic tangent can either saturate towards -1 or 1. Given that the hyperbolic tangent is an asymptotic function near its saturation point, an approximate equality must be considered, as the classification of its saturation will always remain qualitative ($|\alpha_{i}|\neq 1\;\forall\;s_{t}\in B$) . To approximate the condition given in Eq. 2, the amount of dying hyperbolic tangents is calculated by using a kernel density estimation (KDE) (Silverman, 1986) on the activations $\alpha_{i},i\in\mathbb{R}^{w}$ of each individual neuron in the activation layer. In order to visualize activations in a hidden layer, a fixed subset of the KDE’s of the neurons $\alpha_{i}$ is taken. A clear visualization of dying hyperbolic tangents during training in the Atari Breakout environment can be seen by analyzing sixteen individual neuron KDE’s in Fig. 3.
@@
As Fig. 3 indicates that the activation of $z_{t}$ with a hyperbolic tangent leads to saturation and dying neurons, an augmentation of the representation architecture is proposed. In the conventional encoder setting, the networks’ final hidden layer is defined as $z^{enc}(x)=f(A_{1}x+B_{1})$, with $A_{i}$ and $B_{i}$ representing weight and bias parameters, the function $f()$ representing a nonlinear activation function while $x$ is the set of activations from the previous layer. In order to reduce the information dependence on a single set of neurons, we propose using a Hadamard representation that augments the original representation with a parallel representation layer $z^{*}$. This can be interpreted as using a single highway layer with a closed carry gate (Srivastava et al., 2015). The final representation is defined as the Hadamard product between the aforementioned activations $z(x)=z^{enc}(x)\cdot z^{*}(x)$, where $z^{*}(x)=f(A_{2}x+B_{2})$. A visualization of the proposed architecture can be found in Fig. 4.
Figure 4: A visualisation of the Hadamard Representation architecture combined with the nature DQN architecture applied on a snapshot of the ‘SpaceInvaders’ Atari environment (Mnih et al., 2015). Vertical and horizontal bars represent convolutional and linear layers, respectively. A parallel linear layer providing independently parameterized hidden layer activations is integrated, where the Hadamard product of the two activations represents the final latent representation $z_{t}$ used for downstream learning.
@@
$h^{\prime}(x)=\uline{\color[rgb]{0.875,0.26171875,0.8984375}{\color[rgb]{0,0,0}A_{1}\text{sech}^{2}(A_{1}x+B_{1})\tanh(A_{2}x+B_{2})}}\color[rgb]{0,0,0}+\uline{\color[rgb]{0.296875,0.76171875,0.83203125}{\color[rgb]{0,0,0}A_{2}\text{sech}^{2}(A_{2}x+B_{2})\tanh(A_{1}x+B_{1})}}$
where $\operatorname{sech}^{2}$ is the derivative of a hyperbolic tangent. When using a hyperbolic tangent, the key difference with a sigmoid is that for any saturating value of $g(x)$, it will consistently keep the second part of the equation nonzero and dominated by $h(x)$, due to the hyperbolic tangents’ saturation into strictly nonzero values. This means that during saturation of a neuron in $g(x)$, the non-saturated function $h(x)$ is still able to maintain a non-trivial gradient in the product derivative $h^{\prime}(x)$, thus providing a mechanism to avoid vanishing gradients. We visualize the kernel densities during training with a Hadamard representation in Fig. 5. The individual representations before taking the Hadamard product can be found in Appendix B.
Taking a more formal approximation of neuron collapse, we start by defining the probability of a single neuron saturating as $\mathbf{p}$. Furthermore, in the case of a sigmoid or hyperbolic tangent, we assume symmetric saturation probabilities to both ends, defining the probability of a neuron saturating to one end of the spectrum as $0.5\mathbf{p}$. Under these assumptions, we show that interpreting a neuron as the product of two individual neurons will change saturation probabilities depending on the neuron’s activation function.
@@ 4 Experiments
The main experiments and discussion are conducted with the value-based DQN algorithm (Mnih et al., 2015) in Atari, and complemented by an evaluation on the policy-based proximal policy optimization algorithm (PPO) (Schulman et al., 2017). In all experiments, we build on the implementations from cleanrl  (Huang et al., 2022). For more details on the PPO and DQN implementation, we refer the reader to Appendix A.1. For both DQN and PPO, we define the representation $z_{t}\in\mathbb{R}^{512}$ as the last hidden layer of the network, which serves as the compressed representation of the original pixel observation $s_{t}$ (see Fig. 4).
We perform all experiments on 8 common, non-exploration driven Atari games. We aim to show the effect of a Hadamard representation on the representation’s fraction of dying neurons, its effective rank and its downstream performance. For more information on the dying neuron and effective rank calculations, we refer the reader to Appendix. B.1 and Appendix. B.2, respectively. The performance is aggregated over 8 environments and normalized using the ReLU baseline scores.
@@
For more dying neuron graphs related to the ReLU and the Sigmoid activations, we refer the reader to Appendix C.3.
@@
We additionally investigate the effective rank (Kumar et al., 2021) of the representation $z_{t}$ during training, which can be seen in Fig. 6(b). As observed by Gulcehre et al. (2022), a representation activated by a hyperbolic tangent or a sigmoid, already has a relatively high effective rank compared to a representation activated by a ReLU. Furthermore, similar to the results in our supervised learning experiments (see Fig. 2), using a Hadamard representation with hyperbolic tangents significantly improves the effective rank, which is strongly correlated to a network’s ‘expressivity’. Employing a Hadamard representation with ReLU activations significantly decreased the effective rank of the representation, as was expected. For more information on the effect of a Hadamard representation on a ReLU, we refer the reader to Appendix C.4.
@@
The influence of a Hadamard representation on downstream performance can be seen in Fig. 7. Correlating with the reduction in dying neurons and an increase in effective rank, a significant improvement of nearly $60\%$ over the standard hyperbolic tangent baseline is obtained, as well as an improvement over the default ReLU baseline. Note that no hyperparameter or architectural tuning has been employed in our experiments.
@@
Figure 7: Normalized performance with the standard deviation over the means in the Atari domain, after training DQN for 10M iterations (40M Frames). In (a), the Tanh (HR) significantly outperforms a the baseline Tanh. In (b), several ablations including layer normalization (Ba et al., 2016) on hyperbolic tangent activations are shown. Notably, the decrease in performance when using a conventional representation with twice as much neurons ($z_{t}\in\mathbb{R}^{1024}$) shows that more parameters are not always desirable. Additional experiments combining the ReLU activation with a Hadamard representation can be found in Appendix C.4.
@@
Examining further ablations in Fig. 8 shows that using a sophisticated Linear-Unit function such as the SELU activation (Klambauer et al., 2017) or using an addition rather than a product of hyperbolic tangents seems detrimental to performance. Furthermore, taking a product of 3 hyperbolic tangents (2HR) also appears to enhance performance, though there seems to be a negative effect in the early stages of training as compared to using a single Hadamard product. We hypothesize that this is the result of increased contracting behavior in the early stage of training due to increasing multiplication of hyperbolic tangent activations whose absolute values are $<1$. For human-normalized scores, we refer the reader to Appendix D.1.
@@
Interestingly, it seems that the ReLU activations’ performance is less correlated to its low effective rank and high amount of dying neurons than the continuously differentiable activations. As also indicated by Teney et al. (2024), the exact reasons for the broad success of the ReLU activations are not yet fully understood. A qualitative overview of our results can be found in Table 3.
Table 3: High-level summary of the strengths of a Hadamard representation.
Activation function Performance Few dead neurons Effective rank
Tanh (HR) ++ ++ ++
Tanh - - +
ReLU +
Sigmoid - - -
Increasing Representation Dimensionality
A Hadamard representation increases the amount of trainable parameters in the encoder without increasing the dimension of the representation $z_{t}\in\mathbb{R}^{512}$. This is an important feature as simply increasing the dimension of the representation to $\mathbb{R}^{1024}$ is not often effective (see Fig. 7(b)). Even though the fraction of dying neurons decreases and the effective rank increases when using $z_{t}\in\mathbb{R}^{1024}$, it seems that the ’relative’ effective rank, defined as the ratio $\frac{rank}{dim(z_{t})}$, actually goes down (See Appendix. C.2). We therefore hypothesize that reducing dying neurons and increasing a representation’s effective rank is only superficially solved by expanding representation dimensionality.

### v3

@@ Abstract
Activation functions are one of the key components of a deep neural network. The most commonly used activation functions can be classed into the category of continuously differentiable (e.g. tanh) and piece-wise linear functions (e.g. ReLU), both having their own strengths and drawbacks with respect to downstream performance and representation capacity through learning (e.g. measured by the number of dead neurons and the effective rank). In reinforcement learning, the performance of continuously differentiable activations often falls short as compared to piece-wise linear functions. We provide insights into the vanishing gradients associated with the former, and show that the dying neuron problem is not exclusive to ReLU’s. To alleviate vanishing gradients and the resulting dying neuron problem occurring with continuously differentiable activations, we propose a Hadamard representation. Using deep Q-networks, proximal policy optimization and parallelized Q-networks in the Atari domain, we show faster learning, a reduction in dead neurons and increased effective rank.
@@ 1 Introduction
In reinforcement learning (RL) (Sutton & Barto, 2018), the dying neuron phenomenon is more prevalent than in supervised learning due to the use of non-stationary targets (Sokar et al., 2023). However, even though training results in a large number of dying ReLU’s (Gulcehre et al., 2022; Sokar et al., 2023), the ReLU function still remains the most popular activation for performance reasons (Henderson et al., 2018). Similar to supervised learning (Teney et al., 2024), continuously differentiable activation functions such as the hyperbolic tangent are therefore not commonly used in RL (see Fig. 1). However, one might argue that their symmetrical, bounded shape and smooth gradient landscape offer optimization advantages that the ReLU lacks. Recent findings also indicate that a hidden layer activated by a hyperbolic tangent displays a high effective rank and thus a high layer expressivity (Kumar et al., 2021; Gulcehre et al., 2022). Despite being a theoretically sound candidate, its lack of success in RL has not been thoroughly investigated.
We therefore aim to provide insights into the hyperbolic tangent’s suboptimality, revealing that the vanishing gradient problem leads to dying neurons and under-utilization of the full network capacity. Based on these insights, we mitigate said effects by augmenting the original architecture. Specifically, we focus our research on the activations of an encoder, defined as the hidden layer, and provide an alternative to the conventional parameterization of this representation. Our contributions can be summarized as follows:
  • We show that, in reinforcement learning, dying hyperbolic tangents are a phenomenon of a similar scale as the dying ReLU problem, but have a more profound effect on performance.
@@
  • We empirically show that, without hyperparameter tuning or the use of auxiliary losses, using Hadamard representations yields notable performance gains in multiple algorithms in the Atari domain, and reveal how it decreases dying neurons and increases the internal representations’ effective rank.
@@ 3 Augmenting Hyperbolic Tangents
Continuously differentiable activations such as the hyperbolic tangent ($\tanh$) and the sigmoid ($\sigma$) activations are fundamentally different than the ReLU or its piece-wise linear descendants, which are non-symmetric and have a large part of the input space mapped to zero (leading to sparsity). The hyperbolic tangent and the sigmoid output values in the ranges $[-1,1]$ and $[0,1]$, respectively. These functions are defined as $\tanh(x)=\frac{e^{x}-e^{-x}}{e^{x}+e^{-x}}$ and $\sigma(x)=\frac{1}{1+e^{-x}}$.
@@
Although common literature has focused on the dying ReLU problem (He et al., 2015; Lu et al., 2019; Gulcehre et al., 2022; Sokar et al., 2023), we find that hidden layers activated by hyperbolic tangents similarly show strong dying neuron behavior. When using any activation function in a deep neural network, a single neuron $\alpha_{i},i\in\mathbb{R}^{w}$, with $w$ the layer dimension, is saturated or dying if:
$\alpha_{i}\approx\Omega\;\forall\;s_{t}\in B$ (1)
Where $\Omega$ represents the saturation value and $s_{t}$ is an observation in buffer $B$. In practice, a mini-batch of observations is evaluated instead of the whole dataset in $B$. For the hyperbolic tangent, given that it is an asymptotic function near its saturation point, an approximate equality must be considered, as the classification of its saturation will always remain qualitative ($|\alpha_{i}|\neq 1\;\forall\;s_{t}\in B$) . To approximate the condition given in Eq. 1, the amount of dying hyperbolic tangents is calculated by using a kernel density estimation (KDE) (Silverman, 1986) on the activations $\alpha_{i},i\in\mathbb{R}^{w}$ of each individual neuron in the activation layer. In order to visualize activations in a hidden layer, a fixed subset of the KDE’s of the neurons $\alpha_{i}$ is taken. A clear visualization of dying hyperbolic tangents during training in the Atari Breakout environment can be seen by analyzing sixteen individual neuron KDE’s in Fig. 3.
@@
As Fig. 3 indicates that the activation of $z_{t}$ with a hyperbolic tangent leads to saturation and dying neurons, an augmentation of the representation architecture is proposed. In the conventional encoder setting, the networks’ final hidden layer is defined as $z^{enc}(x)=f(A_{1}x+B_{1})$, with $A_{i}$ and $B_{i}$ representing weight and bias parameters, the function $f()$ representing a nonlinear activation function while $x$ is the set of activations from the previous layer. In order to reduce the information dependence on a single set of neurons, we propose using a Hadamard representation that augments the original representation with a parallel representation layer $z^{*}$. This can be interpreted as using a single highway layer with a closed carry gate (Srivastava et al., 2015), or as an augmented version of the Gated Linear Unit (GLU) (Dauphin et al., 2017). The final representation is defined as the Hadamard product between the aforementioned activations $z(x)=z^{enc}(x)\cdot z^{*}(x)$, where $z^{*}(x)=f(A_{2}x+B_{2})$. A visualization of the proposed architecture can be found in Fig. 4.
Figure 4: A visualisation of the Hadamard Representation architecture combined with the nature DQN architecture applied on a snapshot of the ‘SpaceInvaders’ Atari environment (Mnih et al., 2015). Vertical and horizontal bars represent convolutional and linear layers, respectively. A parallel hidden layer providing independently parameterized activations is integrated, where the Hadamard product of the two hidden layers represents the final latent representation $z_{t}$ used for downstream learning.
@@
$z^{\prime}(x)=\uline{\color[rgb]{0.875,0.26171875,0.8984375}{\color[rgb]{0,0,0}A_{1}\text{sech}^{2}(A_{1}x+B_{1})\tanh(A_{2}x+B_{2})}}\color[rgb]{0,0,0}+\uline{\color[rgb]{0.296875,0.76171875,0.83203125}{\color[rgb]{0,0,0}A_{2}\text{sech}^{2}(A_{2}x+B_{2})\tanh(A_{1}x+B_{1})}}$
In this context, $\operatorname{sech}^{2}$ is the derivative of the hyperbolic tangent function. Unlike the sigmoid, the hyperbolic tangent saturates to nonzero values, ensuring that if and only if both parts are saturated, product saturation occours. Thus, when $g(x)$ saturates, $h(x)$ still keeps a non-trivial gradient in the product, providing a mechanism to avoid vanishing gradients. We visualize the kernel densities during training with a Hadamard representation in Fig. 5. The individual representations before taking the Hadamard product can be found in Appendix B.
Taking a more formal approximation of neuron collapse, we start by defining the probability of a single neuron saturating as $\mathbf{p}$. Furthermore, in the case of a sigmoid or hyperbolic tangent, we assume symmetric saturation probabilities to both ends, defining the probability of a neuron saturating to one end of the spectrum as $0.5\mathbf{p}$. Lastly, we make an independence assumption between two individual neurons. Under these assumptions, we show that interpreting a neuron as the product of two individual neurons can change saturation probabilities depending on the neuron’s activation function.
@@ 4 Experiments
The main experiments and discussion are conducted with the value-based DQN algorithm (Mnih et al., 2015) in Atari, and complemented by an evaluation on the policy-based proximal policy optimization algorithm (PPO) (Schulman et al., 2017) and the parallelized Q-network (PQN) (Gallici et al., 2024). For more details on the PPO, DQN and PQN implementation, we refer the reader to Appendix A.1. In line with Gulcehre et al. (2022), for our qualitative experiments, a Hadamard Representation is only applied to the last hidden layer $z_{t}\in\mathbb{R}^{512}$. We perform qualitative experiments on 8 common, non-exploration driven Atari games. We aim to show the effect of a Hadamard representation on the representation’s fraction of dying neurons, its effective rank and its downstream performance. For more information on the dying neuron and effective rank calculations, we refer the reader to Appendix. B.1 and Appendix. B.2, respectively. The performance is aggregated over 8 environments and normalized using the ReLU baseline scores. Finally, the PQN experiments are conducted with an HR on all hidden layers, showing the Median Human-Normalized scores on 51 Atari games.
@@
For more activation-specific dying neuron graphs, we refer the reader to Appendix C.3.
@@
We additionally investigate the effective rank (Kumar et al., 2021) of the representation $z_{t}$ during training, which can be seen in Fig. 6(b). As observed by Gulcehre et al. (2022), a representation activated by a hyperbolic tangent or a sigmoid, already has a relatively high effective rank compared to a representation activated by a ReLU. Furthermore, similar to the results in our supervised learning experiments (see Fig. 2), using a Hadamard representation with hyperbolic tangents significantly improves the effective rank, which is strongly correlated to a network’s ‘expressivity’. Employing a Hadamard representation with ReLU activations significantly decreased the effective rank of the representation, as was expected.
@@
The influence of a Hadamard representation on downstream performance can be seen in Fig. 7. Correlating with the reduction in dying neurons and an increase in effective rank, a significant improvement over the standard hyperbolic tangent baseline is obtained, as well as an improvement over the default ReLU baseline. Furthermore, a comparison is made with the novel Rational (Delfosse et al., 2024) activation function as the activation in the final hidden layer. Fig. 7(a) shows that, although the Rational activation seems to be a stable learnable activation, it remains comparable to the ReLU.
@@
Figure 7: Normalized performance with the standard deviation over the means in the Atari domain, after training DQN for 10M iterations (40M Frames). In (a), the Tanh (HR) significantly outperforms the baseline Tanh. The recently proposed Rational Activation (Delfosse et al., 2024) seems to have comparable performance to the ReLU in the 40M frame setting. In (b), several ablations including layer normalization (Ba et al., 2016) and an increased dimension of $z_{t}\in\mathbb{R}^{1024}$ with a reduced $\alpha=5e^{-5}$ on hyperbolic tangent activations are shown. Note that no hyperparameter changes are done for the Hadamard Representation.
@@
Examining further ablations in Fig. 8 shows that using a sophisticated piece-wise linear function such as the SELU activation (Klambauer et al., 2017) or using an addition rather than a product of hyperbolic tangents seems detrimental to performance. Furthermore, taking a product of 3 hyperbolic tangents (2HR) also appears to enhance performance, though there seems to be a negative effect in the early stages of training as compared to using a single Hadamard product. We hypothesize that this is the result of increased contracting behavior in the early stage of training due to increasing multiplication of hyperbolic tangent activations whose absolute values are $<1$. Additional experiments combining the ReLU activation with a Hadamard representation can be found in Appendix C.4.
@@
Evaluating Hadamard representations on PQN
To more reliably test the performance gains one can achieve using Hadamard Representations, we evaluate their effects on the Parallelized Q-Network (PQN) in the full Atari suite without hard-exploration environments, for a total of 51 games with 5 seeds per game.
Figure 10: Median Human-Normalized scores on PQN in 51 Atari games for 5 seeds over 40M frames. Labels represent MLP / CNN activations. Employing a Hadamard Representation in the Convolutional layers as well as the MLP layer provides a significant performance improvement over the recent PQN baseline, as well as a CReLU (Abbas et al., 2023) implementation of PQN.
PQN has recently emerged as a vectorized-friendly version of DQN, reporting both a speed up from DQN as well as better convergence (Gallici et al., 2024). We run tests implementing the Hadamard Representation in all hidden layers. The median human-normalized scores after training for 40M frames are shown in Fig. 10. We can clearly see a similar trend for the 51-game Atari suite, where using a Hadamard representation for all layers significantly increases the median performance as compared to the baseline PQN and CReLU (Abbas et al., 2023)
Hyperbolic Tangents turn Weights into Biases
Interestingly, it seems that the ReLU activations’ performance is less correlated to its low effective rank and high amount of dying neurons than the continuously differentiable activations. As also indicated by Teney et al. (2024), the exact reasons for the broad success of the ReLU activations are not yet fully understood. However, in the case of dying neurons, we aim to shed some light on this phenomenon.
When dying neurons occur in ReLU-activated layers, it basically prunes these neurons and the associated weights to the next layer. However, in hyperbolic tangent activated layers, dying neurons lead to an unintended phenomenon where weights associated with dead neurons effectively become biases.
Theorem 4.1.
When any set of neurons $\alpha^{j}$ in a hidden layer $z^{j}$ collapses into nonzero values, the output to the next layer effectively changes from ($A^{j}z^{j}+B^{j}$) to ($A_{-*}^{j}z_{i-*}^{j}+B^{j+1}$ + $A_{*}^{j}z_{i*}^{j}$), where $A_{-*}^{j}z_{-*}^{j}$ represent the active neurons multiplied by their corresponding forward-connected weights and $A_{*}^{j}z_{i*}^{j}~{}=~{}B_{*}^{j+1}$ represent the dead neurons, inducing a hidden bias.
Proof.
Take a set of neurons $\alpha^{j}_{i}$ and forward connected weights $\bm{w}^{j}_{\alpha_{i}}$ in layer $z^{j}$. The influence of these neurons on the next layer $z^{j+1}$ is calculated as:
$z^{j+1}=\sum_{i}\alpha^{j}_{i}\bm{w}^{j}_{\alpha_{i}}+B^{j+1}.$ (2)
If the set of neurons dies and collapses into 0 ($\alpha^{j}_{i}=0~{}\forall~{}s\in\mathcal{S}$), which occurs when using ReLU activations, the influence on the next layer becomes 0, representing basic pruning:
$\sum_{i}0\cdot\bm{\bar{w}}^{j}_{\alpha_{i}}=0\quad\forall s\in\mathcal{S}.$ (3)
Where $\bm{\bar{w}}^{j}_{\alpha_{i}}$ is the set of weights connected to the dying neurons. However, for a hyperbolic tangent activation, if the set of neurons saturate into either -1 or 1 ($\alpha^{j}_{i}=\{1,-1\}\forall s\in\mathcal{S}$), the output is
$\sum_{i}\{1,-1\}\cdot\bm{\bar{w}}^{j}_{\alpha_{i}}=B_{*}^{j+1}\quad\forall s\in\mathcal{S}.$ (4)
As a result, the weight vector $\bm{\bar{w}}^{j}_{\alpha_{i}}$ corresponding to the dying activations only influences a bias $B^{j+1}_{*}$ on the resulting hidden layer. Note that the bias $B^{j+1}_{*}$ is constant for any input observation. ∎
This emergent bias can hinder the optimization process by introducing unintended fixed contributions to a networks’ next hidden layer or Q-values, reducing the flexibility of the network’s representations and potentially reducing its performance in RL or supervised learning.